### v1

# TinyAlign: Boosting Lightweight Vision-Language Models by Mitigating Modal Alignment Bottlenecks

Abstract
Lightweight Vision-Language Models (VLMs) are indispensable for resource-constrained applications. The prevailing approach to aligning vision and language models involves freezing both the vision encoder and the language model while training small connector modules. However, this strategy heavily depends on the intrinsic capabilities of the language model, which can be suboptimal for lightweight models with limited representational capacity. In this work, we investigate this alignment bottleneck through the lens of mutual information, demonstrating that the constrained capacity of the language model inherently limits the Effective Mutual Information (EMI) between multimodal inputs and outputs, thereby compromising alignment quality. To address this challenge, we propose TinyAlign, a novel framework inspired by Retrieval-Augmented Generation, which strategically retrieves relevant context from a memory bank to enrich multimodal inputs and enhance their alignment. Extensive empirical evaluations reveal that TinyAlign significantly reduces training loss, accelerates convergence, and enhances task performance. Remarkably, it allows models to achieve baseline-level performance with only 40% of the fine-tuning data, highlighting exceptional data efficiency. Our work thus offers a practical pathway for developing more capable lightweight VLMs while introducing a fresh theoretical lens to better understand and address alignment bottlenecks in constrained multimodal systems.
1 Introduction
The rapid advancements in Large Language Models (LLMs) have catalyzed the development of Vision-Language Models (VLMs), enabling models to excel in complex multimodal reasoning and understanding tasks. Prominent models such as Gemini 2.5 Pro [12], GPT-4V [29], Qwen2.5-VL 72B [2], and PaLI-X [6] have showcased remarkable performance across various benchmarks, setting new standards for multimodal intelligence. However, these models typically involve billions of parameters, resulting in significant computational and storage demands. Such massive requirements make these models impractical for resource-constrained scenarios, such as edge devices or applications with limited computational budgets. This growing demand for efficiency has turned lightweight VLMs into a critical area of research, as they aim to retain strong multimodal capabilities while drastically reducing computational costs and memory footprints, thereby enabling broader applicability.
To achieve this balance between performance and efficiency, most VLMs adopt a modular design where pre-trained vision encoders and language models are frozen, and a small "connector" module is trained to align the two modalities. This approach, also seen in lightweight models such as MiniGPT-4 [46], BLIP-2 [20], and Visual Instruction Tuning [23], is computationally efficient and leverages the strong representational power of pre-trained components. However, while effective, this implicit alignment paradigm might struggles in the context of lightweight VLMs. The limited representational capacity of smaller LLMs significantly constrains their ability to process and align multimodal information, leading to subpar performance on complex tasks. This misalignment becomes a critical bottleneck for lightweight VLMs, as it prevents these models from fully realizing their potential.
To address this issue, we conduct a theoretical analysis of the alignment bottleneck in lightweight VLMs from the perspective of information theory[47, 24]. Specifically, we introduce the concept of Effective Mutual Information (EMI) to quantify the amount of information a model can practically leverage given its capacity constraints. Our analysis reveals that freezing the parameters of a lightweight language model inherently limits the EMI, creating a bottleneck that restricts the model’s ability to align multimodal inputs effectively. This limitation manifests in suboptimal learning dynamics when using standard training objectives, such as cross-entropy loss, and highlights the need for new strategies to enhance the effective flow of information during alignment.
Motivated by this insight, we propose TinyAlign, a novel pre-training and fine-tuning framework explicitly designed to overcome this alignment bottleneck. Inspired by Retrieval-Augmented Generation (RAGs) [19, 14, 15] , TinyAlign introduces a memory bank constructed from multimodal training instances within the dataset, setting it apart from traditional approaches that depend on external knowledge sources. During training, the framework retrieves contextually relevant representations from the memory bank and augments the original visual inputs with enriched multimodal context. By increasing the effective information content available to the model, TinyAlign reduces the inherent learning difficulty posed by the limited capacity of lightweight language models. This design not only enhances alignment but also optimizes the utilization of available training data, addressing the core challenges faced by lightweight VLMs.
We validate the effectiveness of TinyAlign through extensive experiments across a diverse set of lightweight architectures, including Vicuna [8], Phi-2 [13], TinyLLaMA [44], and Qwen2 [36], as well as vision encoders like SigLIP [43] and CLIP [30]. Our results demonstrate that TinyAlign significantly accelerates convergence, reduces alignment losses (Fig. 1(a)), and produces more compact and meaningful feature representations (Fig. 1(b)). Furthermore, TinyAlign exhibits exceptional data efficiency, achieving baseline-level performance with only 40% of the fine-tuning data. Our contribution can be summarized as:
  • We identify a fundamental alignment bottleneck in lightweight VLMs, theoretically analyzing its root cause using principles from information theory. Our analysis reveals that the limited capacity of lightweight LLMs constrains the effective information flow during multimodal alignment.
  • We propose TinyAlign, a novel framework inspired by Retrieval-Augmented Generation, which enriches multimodal inputs by retrieving relevant context from a memory bank. This approach enhances the effective information content accessible to the model, addressing the alignment bottleneck.
  • We conduct extensive experiments to validate the effectiveness of TinyAlign, demonstrating notable improvements in convergence speed, alignment quality, downstream task performance, and data efficiency compared to conventional alignment methods.
2 Related Work
The LLM-Centric Paradigm in VLMs.
The LLM-Centric Paradigm has become a dominant framework in Vision-Language Models, leveraging pre-trained Large Language Models as the core for cross-modal understanding[36, 2, 25, 20, 4, 7]. This approach typically freezes the parameters of both the vision encoder and the LLM [20], while training a lightweight connector module to bridge vision and language [36]. This implicit alignment strategy has achieved notable empirical success, as seen in models like DeepSeek-VL [25] and Qwen2.5-VL [2]. However, the theoretical mechanisms enabling effective cross-modal harmonization remain underexplored, with most research focusing on empirical results rather than systematic analysis. To fill this gap, our work conducts one of the first in-depth theoretical investigations into this paradigm, uncovering its principles and limitations to better understand cross-modal alignment.
Advancements in Lightweight VLMs.
Recent efforts to create lightweight Vision-Language Models (VLMs) have explored various avenues for enhancing efficiency and performance[41, 45, 39, 27, 33]. EfficientVLM [35] introduces a distill-then-prune framework with modal-adaptive pruning to compress large VLMs effectively. TinyLLaVA [45] explores optimal pairings of language models, vision encoders, and connectors for small-scale VLMs, while MobileVLM and its successor v2 emphasize architectural innovations, high-quality data, and advanced training strategies [9, 10], while SmolVLM [27] further explores tokenization strategies.MiniCPM-V [39] presents a series of efficient Multimodal Large Language Models (MLLMs) designed for on-device deployment, achieved by integrating advanced techniques in architecture, pre-training, and alignment. However, these approaches primarily focus on component optimization, model compression, advanced training strategies, or designing for edge deployment, and seldom question whether the widely adopted implicit alignment paradigm is fundamentally suitable for models with limited capacity. In contrast, we provide a principled analysis demonstrating that this paradigm intrinsically induces higher alignment loss for smaller models, thereby limiting their potential for robust visual understanding and cross-modal alignment.
Retrieval-Augmented Models.
Retrieval-Augmented Generation enhances factual accuracy in NLP by integrating external knowledge retrieval with parametric models [19]. Techniques like unsupervised retriever pre-training enable efficient access to large-scale documents during training and inference [14]. RAG now extends to multimodal tasks, with MM-REACT combining language models and vision experts for complex reasoning [37], and Re-ViLM reducing parameters by storing knowledge externally for image-to-text generation [38]. Frameworks like RAVEN and MuRAG apply retrieval for multitask learning and open-domain question answering [31, 5], while models like REVEAL unify memory, retrieval, and generation across diverse sources [15]. These approaches rely on large external memory banks and retriever modules to broaden knowledge for reasoning-heavy tasks [3, 15, 31]. In contrast, TinyAlign addresses the EMI bottleneck in lightweight VLMs by constructing a memory bank from multimodal training instances. During pre-training and fine-tuning, TinyAlign retrieves relevant representations to augment visual input, increasing mutual information between inputs and outputs and overcoming the limitations of compact models.
3 Theoretical Framework: LLM-Centric Alignment and Its Limitations
3.1 Cross-Entropy Loss in LLM-Centric VLM Pre-training
We begin with an analysis of the commonly used LLM-centric paradigm for Vision-Language Models, where the objective is to align visual information with a pre-trained LLM. In this setup, the visual input is denoted as $X_{V}$, the accompanying textual instruction as $X_{I}$, and the target output, often a textual description or answer, as $L$.
The processing pipeline in this paradigm is structured as follows:
  1. 1.
    A frozen Vision Encoder (e.g., a Vision Transformer, ViT) with parameters $\theta_{\text{ViT}}$ processes the visual input $X_{V}$ to extract visual features: $\,Z_{V}=\text{ViT}(X_{V};\theta_{\text{ViT}})$.
  2. 2.
    A trainable Connector module, parameterized by $\theta_{C}^{*}$, transforms the visual features $Z_{V}$ into embeddings $H_{V}$ compatible with the LLM’s input space: $\,H_{V}=\text{Connector}(Z_{V};\theta_{C}^{*})$.
  3. 3.
    Simultaneously, the textual instruction $X_{I}$ is processed by the frozen LLM’s input embedding layer (parameterized by $\theta_{\text{LLM}}$), generating instruction embeddings: $\,H_{I}=\text{LLM}_{\text{embed}}(X_{I};\theta_{\text{LLM}})$.
  4. 4.
    The visual embeddings $H_{V}$ and instruction embeddings $H_{I}$ are concatenated to form the joint input: $\,H_{\text{in}}=[H_{V},H_{I}]$.
  5. 5.
    Finally, the frozen LLM produces an output distribution over the target $L$: $\,P_{\text{model}}(L|H_{\text{in}};\theta_{\text{LLM}})$.
3.2 The LLM Bottleneck: Irreducible Error and Effective Mutual Information
As discussed before, we believe the LLM-centric paradigm inherently suffers from a bottleneck due to its reliance on implicit alignment mechanisms. To understand this limitation, we analyze it from an information-theoretic perspective and show how it constrains the system’s ability to fully utilize the information in the inputs. The starting point of this analysis is the conditional cross-entropy (CE) loss, which measures the discrepancy between the true conditional distribution of the target labels $L$ and the model’s predictions. This loss can be decomposed as follows:
$\mathcal{L}_{\text{CE}}(\theta_{C}^{*}|X_{V},X_{I})$ $=H(P_{\text{true}}(L|X_{V},X_{I}))$
$\quad+D_{\text{KL}}(P_{\text{true}}(L|X_{V},X_{I})||P_{\text{model}}(L|[H_{V}(\theta_{C}^{*}),H_{I}];\theta_{\text{LLM}}))$ (1)
Here, $H(P_{\text{true}}(L|X_{V},X_{I}))$ represents the true conditional entropy of $L$ given the inputs $(X_{V},X_{I})$, which is independent of the model parameters and reflects the irreducible uncertainty in the labels. The second term, the KL divergence, measures the gap between the true conditional distribution and the model’s predictions. Thus, minimizing $\mathcal{L}_{\text{CE}}(\theta_{C}^{*})$ is equivalent to minimizing the KL divergence term, as the conditional entropy is fixed.
At the beginning of training, the Connector parameters $\theta_{C}^{*}$ are randomly initialized, leading to arbitrary and semantically meaningless visual embeddings $H_{V}(\theta_{C}^{*})$ from the perspective of the frozen LLM ($\theta_{\text{LLM}}$). When the LLM processes these unfamiliar embeddings $H_{V}$ alongside $H_{I}$, it produces outputs that deviate significantly from the true target $L$, resulting in a high initial loss. During training, the Connector learns to transform raw visual features $Z_{V}$ into embeddings $H_{V}$ that the frozen LLM can interpret coherently with $H_{I}$ to predict $L$. This process can be viewed as the Connector "translating" visual information into the LLM’s fixed semantic space. However, even with an optimal Connector $\theta_{C}^{\text{opt}}$ that produces the best possible $H_{V}$, the frozen LLM’s representational capacity is inherently limited. Its ability to integrate and reason over the novel visual embeddings $H_{V}$ is constrained by its fixed architecture and pre-trained knowledge. This introduces an irreducible error, quantified as the average KL divergence that cannot be eliminated even with optimal Connector parameters:
$\bar{\epsilon}_{\theta_{\text{LLM}}}=\mathbb{E}_{(X_{V},X_{I})}\left[\min_{\theta_{C}^{*}}D_{\text{KL}}(P_{\text{true}}(L|X_{V},X_{I})||P_{\text{model}}(L|[\text{Connector}(Z_{V};\theta_{C}^{*}),H_{I}];\theta_{\text{LLM}}))\right]\geq 0$
This irreducible error, $\bar{\epsilon}_{\theta_{\text{LLM}}}$, reflects the frozen LLM’s inability to fully align its semantic space with the optimally transformed visual embeddings $H_{V}$. The minimum achievable CE loss for the system, given the frozen parameters $\theta_{\text{ViT}}$ and $\theta_{\text{LLM}}$, can therefore be expressed as:
$\min_{\theta_{C}^{*}}\mathcal{L}_{\text{CE}}(\theta_{C}^{*})$ $=H(L|X_{V},X_{I})+\bar{\epsilon}_{\theta_{\text{LLM}}}$ (2)
Here, $H(L|X_{V},X_{I})$ represents the irreducible uncertainty in the target labels, while $\bar{\epsilon}_{\theta_{\text{LLM}}}$ quantifies the frozen LLM’s limitations. To further analyze this bottleneck, we rewrite the minimum loss using the definition of mutual information $I(X;Y)=H(Y)-H(Y|X)$. Substituting $H(L|X_{V},X_{I})$ into Eq. (2), we obtain:
$\min_{\theta_{C}^{*}}\mathcal{L}_{\text{CE}}(\theta_{C}^{*})=H(L)-I(X_{V},X_{I};L)+\bar{\epsilon}_{\theta_{\text{LLM}}}$ (3)
Here, $H(L)$ is the entropy of the target labels, and $I(X_{V},X_{I};L)$ is the true mutual information between the inputs $(X_{V},X_{I})$ and the labels $L$. To capture the system’s actual capability to utilize information, we define an "Effective Mutual Information" ($I_{\text{eff}}$), which represents the mutual information the system can effectively leverage, accounting for the irreducible error $\bar{\epsilon}_{\theta_{\text{LLM}}}$:
$I_{\text{eff}}(X_{V},X_{I};L|\theta_{\text{LLM}},\theta_{\text{ViT}})=I(X_{V},X_{I};L)-\bar{\epsilon}_{\theta_{\text{LLM}}}$ (4)
Substituting $I_{\text{eff}}$ into Eq. (3), the minimum achievable loss becomes:
$\min_{\theta_{C}^{*}}\mathcal{L}_{\text{CE}}(\theta_{C}^{*})\approx H(L)-I_{\text{eff}}(X_{V},X_{I};L|\theta_{\text{LLM}},\theta_{\text{ViT}})$ (5)
This result highlights a critical limitation of the LLM-centric paradigm: the effective mutual information ($I_{\text{eff}}$) is fundamentally constrained by the frozen LLM’s ability to process and integrate visual embeddings. Even with perfect transformation of visual features by the Connector, the irreducible error $\bar{\epsilon}_{\theta_{\text{LLM}}}$ reduces the system’s ability to fully utilize the true mutual information $I(X_{V},X_{I};L)$. It underscores the need for alternative approaches to achieve better cross-modal alignment.
Hypothesis on LLM Scale. Based on the theoretical analysis above, we hypothesize that a smaller, lightweight LLM (with parameters $\theta_{\text{LLM,small}}$) possesses a more restricted representational capacity and is inherently less capable of interpreting and integrating novel semantic inputs $H_{V}$ compared to a larger, more powerful LLM (with parameters $\theta_{\text{LLM,large}}$). As shown in Fig. 2(a), the larger LLM exhibits a low loss even before ever seeing the image, suggesting its superior capacity to handle semantic information. This implies:
$\bar{\epsilon}_{\theta_{\text{LLM,small}}}\geq\bar{\epsilon}_{\theta_{\text{LLM,large}}}$
Consequently, from Eq. (4), the effective mutual information for a system using a smaller LLM will be lower:
$I_{\text{eff}}(X_{V},X_{I};L|\theta_{\text{LLM,small}},\theta_{\text{ViT}})\leq I_{\text{eff}}(X_{V},X_{I};L|\theta_{\text{LLM,large}},\theta_{\text{ViT}})$
And from Eq. (2), its minimum achievable CE loss will be higher:
$\min_{\theta_{C}^{*}}\mathcal{L}_{\text{CE}}(\text{using }\theta_{\text{LLM,small}})\geq\min_{\theta_{C}^{*}}\mathcal{L}_{\text{CE}}(\text{using }\theta_{\text{LLM,large}})$
This theoretical observation explains why VLMs built upon smaller frozen LLMs often converge to higher loss values and exhibit poorer performance. The LLM’s intrinsic capacity to process and integrate the optimally transformed visual embeddings $H_{V}$ becomes the bottleneck for the entire VLM’s capability.
4 TinyAlign: Mitigating Modal Alignment Bottleneck of Lightweight VLMs
Figure 1: Architectural overview of TinyAlign. Given an input image $X_{V}$ and instruction $X_{I}$, a query key derived from these inputs retrieves $k$ similar, Perceiver-compressed multimodal embeddings $E_{R}=\{E_{R_{j}}\}_{j=1}^{k}$ from a pre-constructed Memory Bank. These cues $E_{R}$ are processed by a trainable RAG Connector ($\theta_{RC}^{*}$) into an auxiliary representation $H_{R}$. Concurrently, $X_{V}$ is processed by a Vision Transformer (ViT) and a primary Connector ($\theta_{C}^{*}$) into visual features $H_{V}$. The instruction $X_{I}$ is embedded as $H_{I}$. Finally, a frozen LLM receives the composite input $H_{\text{in}}^{\prime}=[H_{V},H_{R},H_{I}]$ for prediction. This architecture enhances lightweight VLMs by supplying efficiently processed, relevant contextual information, thereby alleviating the LLM’s processing burden.
4.1 Theoretical Analysis: Enhancing Effective Mutual Information via RAG
Lightweight, frozen LLMs ($\theta_{\text{LLM,small}}$) often exhibit high irreducible error $\bar{\epsilon}_{\theta_{\text{LLM}}}$. TinyAlign (Fig. 1), a Retrieval-Augmented Generation (RAG)-enhanced connector architecture, mitigates this by boosting effective mutual information ($I_{\text{eff}}$) with strategically compressed, relevant contextual cues, thereby reducing the LLM’s intrinsic processing demands.
A standard VLM maps visual input $X_{V}$ (via ViT $\theta_{\text{ViT}}$ and primary connector $\theta_{C}^{*}$) to $H_{V}$, with instruction embeddings $H_{I}$. TinyAlign augments this by: 1) retrieving $k$ pre-compressed embeddings $E_{R}$ from a memory bank $\mathcal{M}$ (Sec. 4.2); 2) employing a trainable RAG connector ($\theta_{RC}^{*}$) to transform $E_{R}$ into supplementary representations $H_{R}$; and 3) presenting a composite input $H_{\text{in}}^{\prime}=[H_{V},H_{R},H_{I}]$ to the frozen LLM. Trainable parameters are $\Theta_{C}^{*}=\{\theta_{C}^{*},\theta_{RC}^{*}\}$. We posit that incorporating $E_{R}$ (forming augmented context $X^{\prime}=(X_{V},E_{R},X_{I})$) enhances $I_{\text{eff}}$. The change, $\Delta I_{\text{eff}}$, is:
$\Delta I_{\text{eff}}$ $=[I(X^{\prime};L)-\bar{\epsilon}_{\theta_{\text{LLM}}}(X^{\prime})]-[I(X_{V},X_{I};L)-\bar{\epsilon}_{\theta_{\text{LLM}}}(X_{V},X_{I})]$
$=\underbrace{I(E_{R};L|X_{V},X_{I})}_{\Delta I_{\text{true}}}+\underbrace{(\bar{\epsilon}_{\theta_{\text{LLM}}}(X_{V},X_{I})-\bar{\epsilon}_{\theta_{\text{LLM}}}(X^{\prime}))}_{\Delta\bar{\epsilon}_{\theta_{\text{LLM}}}}$ (6)
$\Delta I_{\text{true}}>0$ as $E_{R}$ (from pertinent captions) provides novel information about target $L$. $\Delta\bar{\epsilon}_{\theta_{\text{LLM}}}>0$ signifies reduced LLM irreducible error. The RAG connector $\theta_{RC}^{*}$ transforms $E_{R}$ into ’LLM-assimilable contextual hints’ from pre-compressed image-text pairs ($\theta_{P}$-processed), making $H_{\text{in}}^{\prime}$ more ’input-friendly’. This enables the fixed-capacity LLM to better approximate $P_{\text{true}}(L|X^{\prime})$, reducing misinterpretations ($\bar{\epsilon}_{\theta_{\text{LLM}}}(X^{\prime})<\bar{\epsilon}_{\theta_{\text{LLM}}}(X_{V},X_{I})$) and alleviating cognitive load, corroborated by faster convergence (Fig. 2(a)),For more theoretical details on introducing RAG to enhance effective mutual information, please refer to Appendix A.
This $\Delta\bar{\epsilon}_{\theta_{\text{LLM}}}$ reduction is crucial for lightweight VLMs with high baseline error. TinyAlign’s architecture (LLM-independent, Perceiver-based pre-compression and efficient RAG connector) maximizes this reduction, acting as a ’cognitive scaffold’ and lowering the reasoning threshold. Consequently, when $\Delta I_{\text{eff}}>0$ (driven by $\Delta I_{\text{true}}>0$ and substantial $\Delta\bar{\epsilon}_{\theta_{\text{LLM}}}>0$), the minimum achievable CE loss is reduced: $\min_{\theta_{C}^{*}}\mathcal{L}_{\text{CE}}(\text{RAG-enhanced})<\min_{\theta_{C}^{*}}\mathcal{L}_{\text{CE}}(\text{standard})$. RAG-enhanced connectors elevate VLM performance by mitigating the burden on frozen (especially lightweight) LLMs using pre-compressed cues, increasing leverageable information despite fixed LLM capacity.
4.2 Memory Bank Design for Lightweight Efficiency
TinyAlign’s memory bank $\mathcal{M}$ supports lightweight VLMs by storing extensive multimodal information as compact key-value $(K_{m},V_{m})$ pairs from a large-scale pre-training dataset, prioritizing minimal storage and low retrieval latency, and we sample 100K image-text pairs from pre-training dataset declared in Sec. 5.1.
Key Generation.
Each key $K_{m}\in\mathbb{R}^{d_{k}}$ constitutes a compact, low-dimensional embedding derived from a source image-text pair $(X_{V_{m}},X_{I_{m}})$. Its generation employs an attention-based aggregation mechanism. Specifically, multi-modal input features first pass through a self-attention layer to yield an attention matrix, capturing contextual dependencies within the inputs. These original features are then element-wise modulated by this attention matrix. Subsequently, the resultant attention-weighted features are subjected to a mean pooling operation across the token dimension(’Vector transform’ in Fig. 1), producing the distilled key vector $K_{m}$. This process is designed to distill salient cross-modal information into a dense representation, thereby facilitating efficient similarity search and establishing $K_{m}$ as a unified query anchor.
Value Generation.
Each value $V_{m}\in\mathbb{R}^{d_{v}}$ is the corresponding compressed latent multimodal embedding $E_{R_{m}}$. An LLM-independent Perceiver[17] model ($\theta_{P}$, Table 3) pre-processes original multimodal instances $\{(X_{V_{j}},\text{Cap}_{j})\}$ into these values. This pre-compression optimizes for inference speed and reduced memory transfer.
This dual-compression strategy (compact keys $K_{m}$ for rapid search, highly compressed values $V_{m}$ for minimal overhead) is central to TinyAlign’s efficiency, providing potent contextual cues with parsimonious resource use.
4.3 Integrated Pre-training and Instruction Tuning Stages
TinyAlign employs a two-stage training strategy, with both stages using the same objective from Eq. (5): initial connector pre-training with a frozen lightweight LLM and ViT, followed by joint LLM and connector fine-tuning with a frozen ViT for task adaptation.
Pre-training Stage. Connectors $\theta_{C}^{*}$ and $\theta_{RC}^{*}$ are trained while the vision encoder ($\theta_{\text{ViT}}$) and LLM ($\theta_{\text{LLM,small}}$) remain frozen. This compels connectors to optimally format retrieved $E_{R}$ (forming augmented input $H_{\text{in}}^{\prime}=[H_{V},H_{R},H_{I}]$) for the LLM’s extant capabilities, addressing its higher irreducible error and reducing cognitive load. This phase validates the connector’s ability to map the semantic spaces of different pre-trained models to the LLM’s semantic space.
Instruction Tuning Stage. This stage adapts the VLM to downstream tasks. The ViT ($\theta_{\text{ViT}}$) remains frozen. Connectors ($\theta_{C}^{*},\theta_{RC}^{*}$) and, critically, the lightweight LLM ($\theta_{\text{LLM,small}}$) are fine-tuned. With connectors supplying simplified, relevant input, LLM fine-tuning enables specialized reasoning by maximally leveraging these augmented inputs from an active memory bank $\mathcal{M}$. This integrated approach optimizes the LLM’s utilization of the enhanced inputs.
5 Experiments
We conduct a comprehensive suite of experiments to rigorously validate the efficacy of TinyAlign. Our evaluation is designed to: (1) assess its impact on pre-training dynamics, specifically convergence speed and loss reduction (Fig. 2(a)); (2) analyze the quality of the learned representations through UMAP visualizations (Fig. 2(b)); (3) quantify performance improvements on diverse downstream multimodal tasks post-instruction tuning, detailed in Table 1; and (4) measure gains in data efficiency during instruction tuning (Sec. 5.4, Fig. 3), A thorough analysis of data efficiency across various benchmarks can be found in Appendix E (Fig. 4). Additionally, our FLOPs analysis, detailed in Appendix D (Table 4). The extensive ablation studies validating our architectural decisions are presented in Appendix F (Tables 5, 6, and 7). These experiments empirically substantiate our theoretical derivations concerning the enhancement of Effective Mutual Information ($I_{\text{eff}}$).
(a) Pre-training Loss Dynamics
(b) UMAP of LLM Input Space (Phi-2)
Figure 2: (a) Comparison of multimodal pre-training loss on the LLaVA dataset. TinyAlign-enhanced models (dashed lines, "+ RAG") exhibit accelerated convergence and lower final training loss compared to baselines (solid lines) across various model scales (e.g., 16.8% loss reduction for Phi-2.7B, 28.2% for TinyLLaMA-1.1B, 13.9% for Qwen2-0.5B at specified sample points). All models use SigLIP vision encoders, with only connector parameters trained. (b) UMAP visualization of the LLM (for example, Phi-2) input space for (i) baseline(dashed lines), (ii) TinyAlign-enhanced (solid lines), and (iii) a non-aligned model. TinyAlign promotes superior semantic clustering and alignment.
5.1 Experimental Setup
Our framework builds upon lightweight VLMs, drawing inspiration from TinyLLaVA [45]. We primarily utilize the LLaVA dataset family [23] for pre-training analysis and Instrcution tuning, Table 2 provides details on the computational key hyperparameters used for the pre-training and Instruction tuning phases of our experiments. Detailed hyperparameter settings are available in Appendix B.
Datasets.
Pre-training Data: For analyzing pre-training dynamics (Sec. 5.2) and constructing the memory bank (details in Appendix F), we employ the LLaVA pre-training set. This set comprises 558K image-text pairs from a LAION-CC-SBU subset [23], with corresponding annotations. Instruction Tuning Data: For evaluating instruction-following capabilities, we use the LLaVA v1.5 SFT dataset, containing approximately 665K samples. Images are sourced from diverse datasets including COCO [22], GQA [16], TextVQA [32], OCR-VQA [28], and VisualGenome [18].
Models and Baselines.
We validate TinyAlign across several lightweight Large Language Models (LLMs) and vision encoders: Vicuna-7B/13B [8], Phi-2 (2.7B) [13], Gemma (2B) [34], TinyLLaMA (1.1B) [44], and Qwen2-0.5B [2]. For vision encoders, we use SigLIP [43] and CLIP [30]. Consistent with findings in TinyLLaVA [45] that favor SigLIP for lightweight VLMs, we primarily adopt SigLIP. The Connectors utilize a two-layer multilayer perceptron with GELU activation functions to transform the input features.
5.2 Pre-training Performance Analysis
Our empirical results demonstrate TinyAlign’s substantial efficacy in enhancing pre-training for lightweight Vision-Language Models (VLMs). Analysis of training loss trajectories and learned multimodal representations reveals two key benefits.
Accelerated Convergence and Reduced Training Loss. TinyAlign markedly accelerates convergence and reduces final training loss on the LLaVA pre-training dataset. As shown in Fig. 2(a), models incorporating TinyAlign, such as Phi-2 (2.7B), TinyLLaMA (1.1B), and Qwen2-0.5B, achieve significant loss reductions of 16.8%, 28.2%, and 13.9%, respectively, compared to baselines. This improved training efficiency and superior model fit empirically validate our hypothesis (Eq. (2)) that TinyAlign enhances Effective Mutual Information ($I_{\text{eff}}$), thereby lowering the achievable cross-entropy loss ($\mathcal{L}_{\text{CE}}$). These benefits are consistent across models ranging from 0.5B to 2.7B parameters, underscoring TinyAlign’s broad applicability.
Improved Embedding Space Alignment and Semantic Coherence. Qualitatively, TinyAlign cultivates superior multimodal representations. UMAP projections (Fig. 2(b)) reveal that TinyAlign (ii) produces significantly tighter, more separable, and semantically coherent image-text embedding clusters compared to the diffuse embeddings from baselines (i) and the disarray in non-aligned models (iii). This improved structural alignment, indicative of a reduced average embedding error ($\bar{\epsilon}_{\theta_{\text{LLM}}}$), is attributed to contextual cues $H_{R}$ from retrieved information guiding the LLM towards more accurate inter-modal associations. Such well-structured representations are crucial for downstream task performance (see Appendix  C for UMAP details).
In summary, TinyAlign demonstrably improves pre-training efficiency and yields higher-quality multimodal representations. These advancements establish a more robust foundation for downstream instruction tuning and enhanced performance across diverse VLM applications.
5.3 Instruction Tuning Performance Analysis
Post-instruction tuning, TinyAlign variants consistently outperform their respective baselines across a majority of the evaluated benchmarks, as detailed in Table 1. For instance, TinyLLaMA-1.1B equipped with TinyAlign achieves a 74.49% accuracy on VQAv2, marking a 3.51% improvement over its baseline. Similarly, Phi-2-2.7B with TinyAlign attains 78.32% on VQAv2, a 2.93% gain. These improvements generally underscore enhanced generalization capabilities stemming from superior pre-training alignment. However, on the highly complex MM-Vet benchmark, the TinyLLaMA-1.1B variant with TinyAlign exhibits a marginal performance decrease (2.4%). This suggests that for certain exceptionally challenging tasks, particularly with smaller model capacities, the benefits of alignment might be nuanced or necessitate further task-specific tuning.
[…]
Benchmark TinyLLaMA-1.1B Phi-2-2.7B Qwen2-0.5B
Baseline +TinyAlign Baseline +TinyAlign Baseline +TinyAlign
GQA [16] 52.37 56.69 $\uparrow$ 58.44 60.73 $\uparrow$ 56.30 57.58 $\uparrow$
MMMU [42] 29.4 30.2 $\uparrow$ 36.2 37.7 $\uparrow$ 31.0 31.4 $\uparrow$
MM-Vet [40] 26.0 23.6 $\downarrow$ 31.5 34.1 $\uparrow$ 20.9 23.6 $\uparrow$
POPE [21] 84.3 86.0 $\uparrow$ 86.6 88.0 $\uparrow$ 86.4 87.2 $\uparrow$
SQA-I [26] 56.52 59.84 $\uparrow$ 67.28 68.12 $\uparrow$ 59.54 60.24 $\uparrow$
TextVQA [32] 46.33 46.41 $\uparrow$ 50.31 55.48 $\uparrow$ 46.08 46.62 $\uparrow$
VQAV2 [1] 70.98 74.49 $\uparrow$ 75.39 78.32 $\uparrow$ 72.95 74.21 $\uparrow$
MME [11] 1105.10 1200.67 $\uparrow$ 1364.2 1412.1 $\uparrow$ 1170.54 1208.87 $\uparrow$
5.4 Data Efficiency Analysis
TinyAlign demonstrates substantial improvements in data efficiency during instruction tuning, as depicted in Fig. 3. Our analysis reveals that TinyAlign-enhanced models (Phi-2 and TinyLLaMA) can achieve a weighted average accuracy comparable to baselines trained on 100% of the instruction tuning data while utilizing only 40% of the data. This efficiency gain is observed across various capabilities, as the Fig. 4 shows: TinyAlign significantly enhances visual reasoning (e.g., GQA, VQAv2, SQA-I), improves scene text understanding (e.g., TextVQA), and curtails model hallucinations (e.g., POPE), frequently matching or exceeding baseline performance with merely 20-60% of the data. While challenges persist for highly complex tasks such as MMMU, where gains are less uniform, the overall results compellingly underscore TinyAlign’s capacity to facilitate the development of highly capable VLMs under data-constrained scenarios. This reduces reliance on extensive labeled datasets and can accelerate development cycles. A detailed benchmark-specific data efficiency analysis is provided in Appendix E.
Figure 3: Model weighted accuracy, computed per Eq. (8), is evaluated against the percentage of instruction tuning data utilized. TinyAlign models consistently demonstrate superior performance, achieving higher accuracy with substantially less data relative to baseline models.
5.5 Ablation Study
To systematically assess our model architecture, we conduct a series of targeted ablation studies focusing on three key design choices: (1) the size and composition of the RAG knowledge base (Memory Bank) (Table 5), (2) the design methodology for the knowledge base keys (Table 6), and (3) the number of documents (Top-K) retrieved during inference (Table 7).
The findings from these ablations, detailed in Appendix F, inform our final model configuration. We adopt a 100k-entry knowledge base, ensure alignment between the vision encoders used for VLM training and RAG key generation, and retrieve the top-5 documents to augment generation. These choices strike an optimal balance of performance, efficiency, and contextual enrichment.
6 Discussion
Conclusion. Lightweight Vision-Language Models are essential for resource-constrained applications, yet their performance is often limited by alignment bottlenecks caused by the constrained capacity of smaller language models. Through a mutual information perspective, we demonstrate that this limitation reduces the Effective Mutual Information between multimodal inputs and outputs, compromising alignment quality. To address this, we propose TinyAlign, a framework inspired by Retrieval-Augmented Generation, which retrieves relevant context from a memory bank to enrich multimodal inputs. Empirical evaluations show that TinyAlign significantly improves task performance, reduces training loss, and accelerates convergence. Notably, it achieves baseline-level performance with only 40% of the fine-tuning data, offering a highly data-efficient solution for lightweight VLMs. This work provides both a practical framework and a theoretical foundation to address alignment challenges in constrained multimodal systems.
Limitation and Future Work. Though TinyAlign is effective, it has several limitations. Its reliance on a well-designed memory bank may require bespoke construction for different LLMs, and compatibility between the Vision Transformer or querying and memory key generation is essential. Additionally, for challenging tasks or smaller models, further task-specific tuning may be needed. Future work could explore adaptive memory construction and more flexible ViT compatibility to enhance alignment across diverse settings.
Appendix A Detailed Analysis on Enhancing Effective Mutual Information via RAG
As discussed, Lightweight, frozen LLMs ($\theta_{\text{LLM,small}}$) often exhibit performance limitations due to a substantial irreducible error $\bar{\epsilon}_{\theta_{\text{LLM}}}$. To mitigate this, we introduce TinyAlign (Fig. 1), a Retrieval-Augmented Generation (RAG)-enhanced connector architecture. This approach boosts effective mutual information ($I_{\text{eff}}$) by supplying strategically compressed, highly relevant contextual cues, thereby reducing the intrinsic processing demands on the frozen LLM.
A standard VLM processes a visual input $X_{V}$ via a ViT ($\theta_{\text{ViT}}$) to obtain $Z_{V}$, which a primary connector ($\theta_{C}^{*}$) maps to $H_{V}$. Instruction embeddings $H_{I}$ are also generated. TinyAlign augments this by: 1) retrieving $k$ relevant, pre-compressed embeddings $E_{R}=\{E_{R_{j}}\}_{j=1}^{k}$ from a memory bank $\mathcal{M}$ (Sec. 4.2); 2) employing a trainable RAG connector ($\theta_{RC}^{*}$) to transform $E_{R}$ into supplementary representations $H_{R}$; and 3) presenting a composite input $H_{\text{in}}^{\prime}=[H_{V},H_{R},H_{I}]$ to the frozen LLM. The trainable parameters are thus $\Theta_{C}^{*}=\{\theta_{C}^{*},\theta_{RC}^{*}\}$. We posit that incorporating $E_{R}$—forming an augmented context $X^{\prime}=(X_{V},X_{I},E_{R})$—enhances $I_{\text{eff}}(X^{\prime};L|\theta_{\text{LLM}},\theta_{\text{ViT}})$. The change, $\Delta I_{\text{eff}}$, is:
$\Delta I_{\text{eff}}$ $=[I(X^{\prime};L)-\bar{\epsilon}_{\theta_{\text{LLM}}}(X^{\prime})]-[I(X_{V},X_{I};L)-\bar{\epsilon}_{\theta_{\text{LLM}}}(X_{V},X_{I})]$
$=\underbrace{I(E_{R};L|X_{V},X_{I})}_{\Delta I_{\text{true}}}+\underbrace{(\bar{\epsilon}_{\theta_{\text{LLM}}}(X_{V},X_{I})-\bar{\epsilon}_{\theta_{\text{LLM}}}(X^{\prime}))}_{\Delta\bar{\epsilon}_{\theta_{\text{LLM}}}}$ (7)
The first term, $\Delta I_{\text{true}}$, is positive because $E_{R}$, derived from pertinent captions (i.e., clues for the target $L$), provides novel information about $L$ conditioned on $X_{V}$ and $X_{I}$. The second term, $\Delta\bar{\epsilon}_{\theta_{\text{LLM}}}$, signifies a positive reduction in the LLM’s irreducible error. The RAG connector $\theta_{RC}^{*}$ transforms $E_{R}$ into $H_{R}$, providing what we term ’concise, LLM-assimilable contextual hints.’ These hints—originating from pre-compressed image-text pairs (potent clues processed by $\theta_{P}$)—present information in a format more attuned to the LLM’s textual processing strengths than deciphering complex visual semantics solely from $H_{V}$. This enhanced ’input friendliness’ of the augmented input $H_{\text{in}}^{\prime}$ enables the fixed-capacity LLM to approximate the target distribution $P_{\text{true}}(L|X^{\prime})$ with greater fidelity than it could from $(X_{V},X_{I})$ alone, a phenomenon corroborated by faster convergence (Fig. 2(a)). This improved fidelity directly translates into fewer fundamental misinterpretations, thereby lowering the irreducible error: $\bar{\epsilon}_{\theta_{\text{LLM}}}(X^{\prime})<\bar{\epsilon}_{\theta_{\text{LLM}}}(X_{V},X_{I})$. The trainable RAG connector $\theta_{RC}^{*}$ is optimized to generate $H_{R}$ that maximizes this error reduction by effectively complementing $H_{V}$ and $H_{I}$, thus alleviating the LLM’s cognitive burden.
This reduction in irreducible error is especially impactful for lightweight VLMs, whose inherent capacity constraints often yield a higher baseline $\bar{\epsilon}_{\theta_{\text{LLM}}}(X_{V},X_{I})$. TinyAlign’s architecture—emphasizing LLM-independent, Perceiver-based pre-compression for the memory bank and an efficient RAG connector—is engineered to maximize this $\Delta\bar{\epsilon}_{\theta_{\text{LLM}}}$. It functions as a ’cognitive scaffold,’ substantially lowering the reasoning threshold for these less capable LLMs—a benefit less critical for LLMs endowed with robust intrinsic reasoning. Consequently, when $\Delta I_{\text{eff}}>0$—driven by both $\Delta I_{\text{true}}>0$ and a substantial $\Delta\bar{\epsilon}_{\theta_{\text{LLM}}}>0$ for lightweight LLMs—the system’s minimum achievable CE loss is reduced: $\min_{\theta_{C}^{*}}\mathcal{L}_{\text{CE}}(\text{RAG-enhanced})<\min_{\theta_{C}^{*}}\mathcal{L}_{\text{CE}}(\text{standard})$. This framework demonstrates that RAG-enhanced connectors, by strategically mitigating the processing burden and inherent error of a frozen (especially lightweight) LLM using pre-compressed multimodal cues, can elevate VLM performance. This effectively increases the information the LLM can leverage despite its fixed capacity.
Appendix B Hyperparameter Summary
This subsection provides a comprehensive overview of the critical hyperparameters employed throughout our experimental phases, encompassing both pre-training and fine-tuning stages for our proposed models, as well as the specific configuration for the Perceiver model utilized in TinyAlign’s memory bank construction. Table 2 delineates the comparative settings for pre-training and fine-tuning, covering aspects such as batch sizes, learning rates, and optimization strategies. Complementing this, Table 3 itemizes the architectural hyperparameters of the Perceiver model, detailing its latent space dimensions, attention mechanisms, and input specifications. All experiments were conducted on a system equipped with eight NVIDIA RTX 4090 GPUs, each with 48GB of VRAM, ensuring robust and scalable computational support.
Table 2: Key hyperparameters for pre-training and fine-tuning.
Hyperparameter Pre-training Value Fine-tuning Value
Global Batch Size 256 128
Per-device Batch Size (script) 16 12
Gradient Accumulation (script) 1 2
Learning Rate 1e-3 5e-8
LR Scheduler Cosine Cosine
Warmup Ratio 0.03 0.03
Precision FP16 FP16
Optimizer AdamW AdamW
LLM Tuning Frozen Full
Vision Tower Tuning Frozen Frozen
Connector Tuning Full Full
Table 3: PerceiverConfig Hyperparameters
Parameter Value
num_latents 32
d_latents 96
d_model 128
num_self_attends_per_block 8
num_blocks 1
num_self_attention_heads 8
num_cross_attention_heads 8
qk_channels 96
v_channels 96
image_size 384
vocab_size 30522
max_position_embeddings 512
Appendix C UMAP Visualization of Feature Embeddings
To elucidate the latent structure and interrelations between image-derived and text-derived features, we employed Uniform Manifold Approximation and Projection (UMAP) for dimensionality reduction and visualization.
UMAP Fundamentals: UMAP is a non-linear dimensionality reduction algorithm adept at preserving both local and global structures of high-dimensional data within a lower-dimensional embedding. It achieves this by first constructing a high-dimensional graph representation of the data, where edge weights denote the likelihood of connectivity between points. Subsequently, it optimizes a low-dimensional graph to maximize structural similarity to its high-dimensional counterpart. This methodology is rooted in Riemannian geometry and algebraic topology, aiming to model the data’s underlying manifold.
Visualization Methodology: Our UMAP-based visualization procedure comprised the following steps:
  1. 1.
    Input Data Preparation: The UMAP process utilized two sets of high-dimensional feature vectors:
    • Connector Features: Vectors derived from images post-processing by the vision tower and the model’s connector module. These represent visual information conditioned for the language model.
    • LLM Input Embeddings: Corresponding vectors representing embeddings of textual inputs (e.g., questions, prompts) from the language model’s input embedding layer.
    These two feature sets were concatenated. Origin labels (connector feature or LLM embedding) were retained for subsequent differentiated visualization.
  2. 2.
    UMAP Dimensionality Reduction: The UMAP algorithm was applied to the concatenated high-dimensional dataset, configured to reduce feature dimensionality to two. This 2D representation facilitates direct scatter plot visualization. The algorithm learns a mapping that optimally preserves the topological structure of the original feature space in this lower dimension.
  3. 3.
    Output and Visualization: The UMAP process yielded:
    • 2D Coordinates: A two-dimensional coordinate (x, y) for each input feature vector (both connector and LLM embedding).
    • Scatter Plot Visualization: These 2D coordinates were used to generate a scatter plot. Points corresponding to connector features and LLM input embeddings were rendered in distinct colors. This visualization enables qualitative assessment of:
      • The distinctness or overlap between the two feature spaces.
      • The presence of intra-feature-type clusters.
      • The overall geometric relationship between the model’s latent representations of visual and textual information.
Appendix D FLOPs Analysis
This section presents a comparative analysis of the Floating Point Operations (FLOPs) for models based on the Phi-2 architecture, focusing on the computational demands during different stages. As detailed in Table 4, we evaluate both our proposed TinyAlign approach and a baseline model across pre-training and fine-tuning phases. The analysis indicates that while the introduction of RAG tokens and an additional RAG connector in TinyAlign incurs a marginal increase in computational operations, the overall impact on inference efficiency and training throughput remains negligible, underscoring the computational viability of our method.
Table 4: FLOPs analysis for Phi-2 based models. While the introduction of RAG tokens and an additional RAG connector incurs a marginal increase in computational operations, the impact on overall inference efficiency is negligible.
Model Stage Train Steps/Second FLOPs
TinyAlign-finetune 0.075 6.15e18
TinyAlign-pretrain 0.112 2.45e14
Baseline-finetune 0.083 6.12e18
Baseline-pretrain 0.173 2.05e14
Appendix E Detailed Data Efficiency Analysis
TinyAlign models exhibit remarkable data efficiency during instruction tuning, as illustrated comparatively in Fig. 3 (main text) and further detailed per benchmark in Fig. 4. Our comprehensive analysis indicates that TinyAlign-enhanced models, specifically LightweightLLM-Phi-2 and TinyLLaMA, can match the weighted average accuracy of baseline models trained with 100% of the instruction tuning data while using only 40% of this data. A granular examination of individual benchmarks further substantiates TinyAlign’s effectiveness in learning from limited labeled data and enhancing critical VLM capabilities.
Enhanced Visual Reasoning with Greater Efficiency.
On visual reasoning benchmarks, TinyAlign yields substantial data efficiency gains. For GQA, TinyAlign (SigLIP-384-phi-2, 2.7B) with 60% data (59.7 accuracy) surpasses the baseline with 100% data (58.4 accuracy). Similarly, TinyAlign (SigLIP-384-TinyLlama, 1.1B) with just 40% data (52.78 accuracy) outperforms its baseline counterpart using 100% data (52.37 accuracy). On VQAv2, TinyAlign (SigLIP-384-phi-2, 2.7B) at 40% data (75.64 accuracy) matches the baseline’s full-data performance (75.39 accuracy), while TinyAlign (SigLIP-384-TinyLlama, 1.1B) with 40% data (72.50 accuracy) significantly exceeds its baseline with 100% data (70.98 accuracy). For SQA-I, the TinyAlign-enhanced TinyLlama model demonstrates a particularly notable improvement in its performance trajectory with increasing data efficiency, compared to a relatively stagnant baseline.
Improved Scene Text Understanding.
TinyAlign’s advantages extend to scene text understanding. On TextVQA, TinyAlign (SigLIP-384-phi-2, 2.7B) utilizing only 40% of the training data (51.10 accuracy) achieves higher accuracy than the baseline model trained with 100% data (50.31 accuracy), suggesting an improved ability to integrate visual and textual cues from sparser signals.
Effective Suppression of Model Hallucinations.
TinyAlign significantly enhances model factuality by reducing hallucinations, as evidenced on the POPE benchmark. TinyAlign models consistently achieve higher accuracy (implying fewer hallucinations) across various data proportions. For instance, TinyAlign (SigLIP-384-TinyLlama, 1.1B) with only 20% data (85.6 accuracy) surpasses the baseline TinyLlama trained with 100% data (84.3 accuracy). TinyAlign (SigLIP-384-phi-2, 2.7B) exhibits similar data efficiency, reaching the baseline’s peak performance with only 40% of the data.
Broad Applicability and Efficacy on Lightweight Models.
On comprehensive benchmarks such as MME, TinyAlign generally provides consistent performance improvements for both Phi-2 and TinyLlama backbones. For instance, TinyAlign-enhanced models show a clear advantage over baselines across most data regimes on MME. On MM-Vet, data efficiency is also prominent: TinyAlign (SigLIP-384-phi-2, 2.7B) with 20% data (32.2 accuracy) outperforms the baseline with 100% data (31.5 accuracy), while TinyAlign (SigLIP-384-TinyLlama, 1.1B) with 40% data (24.6 accuracy) performs comparably to its baseline with 100% data (24.8 accuracy).
Addressing Challenges in Complex Tasks.
While TinyAlign demonstrates broad advantages, its impact can vary with task complexity. On the highly intricate MMMU benchmark, performance gains from TinyAlign are less consistent across data proportions compared to other benchmarks. Furthermore, on MM-Vet, the baseline TinyLlama model slightly outperforms its TinyAlign counterpart when both are trained with 100% data. These instances suggest avenues for future research, such as optimizing TinyAlign for extremely complex tasks or specific model-data interactions.
To comprehensively evaluate the model’s overall performance across multiple distinct benchmark datasets, we employed weighted average accuracy. This method assigns a corresponding weight based on the sample size (or representative importance) $C_{i}$ of each test set. For the accuracies $Acc_{i}$ obtained for each test set under specific conditions (e.g., a certain training data percentage), the weighted average accuracy $A_{\text{weighted}}$ is calculated as follows:
$A_{\text{weighted}}=\frac{\sum(Acc_{i}\times C_{i})}{\sum C_{i}}$ (8)
This approach allows us to more accurately reflect the model’s combined performance on datasets of varying scales and characteristics, avoiding potential biases from simple averaging.
Figure 4: Detailed data efficiency analysis across individual benchmarks. Performance of TinyAlign-enhanced models (variants of Phi-2 and TinyLLaMA) is compared against their respective baselines at varying percentages of instruction tuning data.
Appendix F Detailed Ablation Studies
We conduct several ablation studies to validate key design choices within the TinyAlign framework. These studies focus on knowledge base (KB) size, vision encoder alignment for RAG key generation, and the number of retrieved documents (Top-K). All ablations use the Phi-2 (2.7B) model as the LLM backbone unless otherwise specified.
Knowledge Base Size.
We evaluate KB sizes of 100k, 300k, and 500k entries, with retrieval fixed at top-10 documents (Table 5). The 100k KB shows comparable, and in some cases superior, performance to larger KBs while offering greater efficiency. This finding suggests that a moderately sized, high-quality KB is sufficient and eliminates the need for excessively large memory banks.
Table 5: Ablation on knowledge base (KB) size for Phi-2 (2.7B) with top-10 retrieval. Performance reported on various benchmarks.
KB Size GQA MMMU MM-Vet POPE SQA-I TextVQA VQAv2 MME
100k 60.73 37.7 34.1 0.880 68.12 55.48 78.32 1412.1
300k 61.27 36.8 31.7 0.871 70.25 55.42 78.58 1407.34
500k 61.38 36.0 31.6 0.875 68.91 56.43 78.57 1401.06
Vision Encoder Alignment for RAG Keys.
We investigate the importance of aligning the vision encoder used for VLM pre-training with the encoder used for generating RAG keys (Table 6). Using stablelm-2-zephyr-1.6b as the VLM, we compare a matched configuration (VLM pre-trained with SigLIP features, RAG keys from SigLIP) against a baseline VLM (pre-trained with CLIP features, no RAG) and a mismatched RAG configuration (VLM pre-trained with CLIP features, RAG keys from SigLIP). The results show significant performance degradation, including a "collapse" on benchmarks like VQAv2, when encoders are mismatched. This highlights the critical importance of feature space consistency between the VLM’s learned representations and the RAG system’s retrieval mechanism. The "TinyAlign (Matched)" configuration, representing our proposed approach, consistently performs on par with or better than the CLIP VLM baseline, demonstrating the advantages of properly aligned RAG systems.
Table 6: Ablation on vision encoder alignment. VLM: stablelm-2-zephyr-1.6b. KB: 100k SigLIP-keyed. "Matched": VLM (SigLIP features) + RAG (SigLIP keys). "Baseline": VLM (CLIP features), no RAG. "Mismatched": VLM (CLIP features) + RAG (SigLIP keys).
Benchmark TinyAlign (Matched: SigLIP VLM + SigLIP RAG) Baseline (CLIP VLM) TinyAlign (Mismatched: CLIP VLM + SigLIP RAG)
GQA 51.28 52.73 40.96
MMMU 0.304 0.313 0.300
MM-Vet 28.4 28.4 24.6
POPE 0.847 0.836 0.783
SQA-I 62.32 62.47 56.17
TextVQA 49.39 46.75 36.71
VQAv2 70.12 69.82 29.62
MME 1218.40 1206.49 1054.78
Table 7: Ablation on top-K retrieval for Phi-2 (2.7B) with a 100k knowledge base.
Top-K GQA MMMU MM-Vet POPE SQA-I TextVQA VQAv2 MME
Top-1 60.6 36.0 33.2 0.874 70.35 55.66 78.21 1411.12
Top-5 61.27 38.1 33.6 0.874 70.00 55.69 78.48 1415.22
Top-10 60.73 37.7 34.1 0.880 68.12 55.48 78.32 1412.1
Top-K Retrieval.
Using the 100k KB, we ablate the number of retrieved documents (Top-K) for Phi-2, comparing top-1, top-5, and top-10 retrieval (Table 7). Retrieving the top-5 documents achieves the best overall balance of performance across benchmarks, providing richer contextual signals than top-1 while avoiding the potential noise or diminishing returns observed with top-10 in some cases.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification:: Yes. Our main contributions are also detailed in Sec. 1 . Also see Sec. 345 and Appendix  A,  E for more theoretical and experimental evidence.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Yes, please see Sec. 6 for limitations. We also reporte the FLOPs in Appendix D.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We detail the assumption and proof of theoretical result in Sec. 3,  4 and Appendix  A.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We use publicly-accessable dataset LLaVA Dataset  [23] and detail the model architecture in Sec. 4 and Fig. 1 . Also,We have provided a detailed description of all the hyperparameters and model components used in the experiment in Appendix B and Sec. 5.1.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We use publicly-accessable dataset LLaVA Dataset  [23].Once the blind review period is finished, we’ll open-source all codes.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Yes, please see Appendix B and Sec. 5.1,  4.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Due to the resource limitation, we do not report error bars. Please note that in Sec. 4, based on the theory, we propose to verify this idea, we designed sets of experiments in Sec. 5 to confirm its applicability.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report the training devices in Appendix B and FLOPs in Table 4.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We followed the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [No]
  49. Justification: This work focuses on academic, publicly-available benchmarks as detailed in Sec. 5. This work is not related to any private or personal data, and there’s no explicit negative social impacts.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [No]
  54. Justification: We do not foresee any high risk for misuse of this work.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Yes, we credited them in appropriate ways.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: Yes, please see Sec.  4, Sec. 5 and Appendix C.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.